# Visualization Tasks for Unlabelled Graphs

### v1

@@ 7.1 Utility
Taxonomies can be judged according to their descriptive, evaluative, and generative power [8]. We provide evidence of the descriptive power of our taxonomy by connecting the fully abstracted task triplets to concrete tasks proposed in the previous work. Evaluative power was one of our major motivations in creating it: we wanted to help future researchers in designing user studies, when they are deciding on what tasks would be appropriate. Our preliminary assessment of multiple visual encoding idioms with the taxonomy does provide some evidence of its evaluative power. We also think the taxonomy has great promise in terms of generative power. We have generated some examples of concrete tasks not previously articulated in task taxonomies considering the combinatorial possibilities of abstract triplets. Moreover, Fig. 5 strongly indicates the potential benefit of future technique-oriented work to address some of the regions where existing methods fall short.
@@ 7.2 Limitations and Future Work
The most obvious limitation of our work is that our assessment of visual encoding idioms is highly subjective, and should be considered preliminary conjectures rather than established facts. We intend our analysis to illustrate the evaluative power of our task taxonomy, not to be the final word. In our assessment, we primarily considered whether the visual encodings make the particular target visually salient, or conversely whether the visually salient features in the encodings correspond to the targets of interest. We drew upon the knowledge of all authors, but in many cases we were well aware that more research is needed to further explore these questions—especially for the less well studied invariant encodings. We call for future empirical work through human-subjects studies that could serve to either confirm or modify these conjectures.
@@ 8 Conclusion
We propose a task taxonomy for unlabelled graph tasks organized by Scope + Action + Target triplets. We validate it by considering its descriptive, generative, and evaluative power. We also consider 6 visual encodings (2 traditional and 4 based on invariant descriptors) and perform a preliminary assessment of their usability with our task taxonomy. We encourage formal human-subject studies using the taxonomy to verify the utility of invariant visual encodings that lack quantitative evaluations, such as BMatrix Network Portrait, Graph Thumbnails, Hop-Census, and Census-Census plots.
Supplemental Materials Index
Supplemental materials are available at https://osf.io/e23mr/. We provide a supplemental PDF with three sections: An extended version of our Fig. 1 literature review table (Supp. §1); a more detailed rationale of the reasoning behind the assessment summarized in Fig. 5 (Supp. §2); and a greatly expanded version of Fig. 4 showing the material we used for preliminary visual encoding evaluation reference, computed from an eclectic collection of 81 unlabelled graphs we contributed in our own previous work [47] (Supp. §3). We also provide spreadsheet versions of Supp. §1 and Supp. §2 as a ZIP file.
Acknowledgements.
The development of this paper took place between October 2024 and April 2025 at the UBC Point Grey campus and in the City of Vancouver, which sit on the traditional, ancestral, unceded territory of the Musqueam, Squamish, and Tsleil-Waututh First Nations. This work was supported in part by NSERC DG RGPIN-2024-06401. We thank Steve Kasica, Firas Moosvi, Francis Nguyen, and Mara Solen for their manuscript reviews.

### v2

@@ 8.1 Utility
Taxonomies can be judged according to their descriptive, evaluative, and generative power [7]. We provide evidence of the descriptive power of our taxonomy by connecting the fully abstracted task triplets to concrete tasks proposed in the previous work. Evaluative power was one of our major motivations in creating it: we wanted to help future researchers in designing user studies, when they are deciding on what tasks would be appropriate. Our preliminary assessment of multiple visual encoding idioms with the taxonomy does provide some evidence of its evaluative and descriptive power. We also think the taxonomy has great promise in terms of generative power. We have generated some examples of concrete tasks not previously articulated in task taxonomies considering the combinatorial possibilities of abstract triplets.
We developed this taxonomy for the pragmatic reason that existing invariant plot methods cannot be evaluated without it. However, conversely, we hope that the existence of this task taxonomy will spur future work. In particular, Fig. 5 strongly indicates the potential benefit of future technique-oriented work to address some of the regions where existing methods fall short.
@@ 8.2 Limitations and Future Work
The most obvious limitation of our work is that our assessment of visual encoding idioms is highly subjective and should be considered an educated guess. We intend our analysis to illustrate the evaluative power of our task taxonomy, where we primarily consider whether the visual encodings make the particular target visually salient, or conversely whether the visually salient features in the encodings correspond to the targets of interest. We drew upon the knowledge of all authors, but in many cases we were well aware that more research is needed to further explore these questions; especially for the less well studied invariant plots. We call for future empirical work through human-subjects studies that could serve to either confirm or modify these conjectures.
@@
We developed our data abstraction (Sec. 3) as a filtering tool to analyze tasks. Future research may also use this model to investigate graph-theoretic and algorithmic properties of unlabelled graphs in relation to semantic richness, with applications that can extend beyond visualization given the ubiquity of data with connected topology [29].
@@ 9 Conclusion
We propose a task taxonomy for unlabelled graph tasks organized by Scope + Action + Target triplets. We validate the descriptive, generative, and evaluative power of this taxonomy by considering 6 visual encodings (2 traditional views and 4 invariant plots) across one task sampled from each triplet. We then perform an educated guess assessment of encoding usability for each task, as an example of how user study design could make use of our taxonomy. We encourage formal human-subject studies using the taxonomy to verify the utility of invariant visual encodings that lack quantitative and qualitative evaluations, such as BMatrix Network Portrait, Graph Thumbnails, Hop-Census, and Census-Census plots.
Acknowledgements
The development of this paper took place between October 2024 and October 2025 at the UBC Point Grey campus and in the City of Vancouver, which sit on the traditional, ancestral, unceded territory of the Musqueam, Squamish, and Tsleil-Waututh First Nations. This work was supported in part by NSERC DG RGPIN-2024-06401. We thank Steve Kasica, Firas Moosvi, Francis Nguyen, and Mara Solen for their manuscript reviews.